### v3

@@ Abstract
This paper first proposes the Halfway Escape Optimization (HEO) algorithm, a novel quantum-inspired metaheuristic designed to address complex optimization problems characterized by rugged landscapes and high-dimensionality with an efficient convergence rate. The study presents a comprehensive comparative evaluation of HEO’s performance against established optimization algorithms, including Particle Swarm Optimization (PSO), Genetic Algorithm (GA), Artificial Fish Swarm Algorithm (AFSA), Grey Wolf Optimizer (GWO), and Quantum behaved Particle Swarm Optimization (QPSO). The primary analysis encompasses 14 benchmark functions with dimension 30, demonstrating HEO’s effectiveness and adaptability in navigating complex optimization landscapes and providing valuable insights into its performance. The simple test of HEO in Traveling Salesman Problem (TSP), Pressure Vessel Design and Tubular Column Design infers its feasibility and potential weakness in real-time applications.
@@
we propose a new algorithm called HEO we provide a comprehensive benchmark of 14 Objective Functions for our HEO algorithm we further evaluate the effectiveness of the new algorithm with 3 classical engineering problems
@@ 1 Introduction
The main motivation behind the development of the Halfway Escape Optimization (HEO) algorithm stems from the limitations of existing optimization methods in terms of efficiency and adaptability to various single-objective optimization problems. HEO is a novel quantum-inspired metaheuristic designed to tackle the complexities of rugged landscapes, high-dimensionality, and multimodal search spaces with a focus on achieving fast convergence rates and reducing the time cost in searching. Inspired by the behavior of quantum particles and the concept of halfway escape, the algorithm offers a versatile and adaptive approach to exploration and exploitation in challenging optimization domains.
This paper provides a comprehensive analysis and evaluation of the HEO algorithm, focusing on its core principles, adaptability, and robustness in navigating diverse optimization landscapes. The unique energy-driven behavior, vibration strategies, and exploratory mechanisms of HEO are examined in the context of solving a range of complex benchmark functions. The evaluations demonstrate HEO’s effectiveness in achieving convergence, balancing exploration and exploitation, and discovering high-quality solutions across a variety of challenging optimization problems.
Comparative studies with other swarm optimization algorithms highlight HEO’s adaptability and robustness in addressing the complexities of multimodal and high-dimensional search spaces, showcasing its fast convergence speed. The results of this study underscore the significance of the HEO algorithm as a promising approach for addressing complex optimization challenges, with implications for a wide range of practical applications including structural design[8][9], path planning[26], circuit planning[23], or model optimization[28].
The experimentation in this paper extends to the classical Traveling Salesman Problem (TSP), where the performance of the HEO algorithm is compared with that of Tabu Search and Random Search. TSP involves finding the shortest possible route for a salesperson to visit a set of cities once and return to the starting city. Because TSP is a well-known NP-Hard Problem in the optimization realm, hence it is a well-testing problem with its exponentially increased time complexity[15]. The comprehensive analysis presented demonstrates the potential of the HEO algorithm as a powerful and adaptive optimization method. Overall, the paper first proposes HEO method and primary testing its effectiveness and convergence in 14 single objective functions.
@@ 2 Related Work
In the realm of optimization, metaheuristic algorithms play a pivotal role in addressing complex and challenging optimization problems. These algorithms draw inspiration from natural phenomena and collective behaviors to navigate high-dimensional and multimodal search spaces, offering adaptive and robust solutions.
@@
Particle Swarm Optimization (PSO) has gained widespread recognition as a bio-inspired metaheuristic algorithm rooted in the collective behavior of bird flocks. PSO was first proposed by Eberhart and Kennedy in 1995. PSO iteratively updates particle positions and velocities based on their individual and global best solutions, enabling effective exploration and exploitation of the search space[7]. While PSO has exhibited robustness in addressing various optimization problems, its potential for premature convergence and challenges in handling complex multimodal landscapes have been noted[27].
Ant Colony Optimization (ACO) is an algorithm inspired by the foraging behavior of ants. It was initially proposed by Marco Dorigo in his doctoral thesis in 1992 and has since gained significant attention in the field of optimization[5]. ACO simulates the behavior of ants in finding the shortest path between their nest and food source by depositing pheromone trails. These trails act as a form of communication, allowing ants to navigate and explore the search space effectively. ACO has been successfully applied to a wide range of optimization problems, including the well-known Traveling Salesman Problem (TSP) and Vehicle Routing Problem (VRP)[5]. Its ability to find near-optimal solutions and its adaptability to dynamic environments make ACO a valuable algorithm in solving complex optimization problems.
Differential Evolution (DE) is an another population-based stochastic optimization algorithm that was introduced by [25]. DE operates by iteratively evolving a population of candidate solutions through a combination of mutation, crossover, and selection operations. It is inspired by the process of natural evolution and survival of the fittest. DE utilizes the differential operators to explore and exploit the search space efficiently. It has been widely applied to various optimization problems and has shown promising performance in solving complex and multimodal functions. DE’s simplicity, effectiveness, and robustness make it a popular choice for solving real-world optimization problems.
@@
These existing metaheuristic algorithms have significantly contributed to the field of optimization, each with its unique strengths and limitations, the development of the Halfway Escape Optimization (HEO) algorithm seeks to offer a novel and adaptive approach to address the challenges of rugged landscapes, high-dimensionality, and multimodal search spaces [27].
To overcome the challenges presented by rugged landscapes, high dimensionality, and multimodal search spaces, a novel approach is required. Yet QPSO and HEO are both quantum-inspired algorithms with the same mechanism in computing global weight for updating position(seen in equation (6)), they still have significant differences. In contrast to the Quantum Particle Swarm Optimization (QPSO), the Halfway Escape Optimization (HEO) algorithm aims to provide such an approach by leveraging insights from quantum patterns and halfway escape strategies.
@@ 3 Proposed Methodology
In recent years, swarm optimization algorithms have been widely applied in various fields, such as machine learning, engineering design. However, these algorithms often encounter difficulties when facing complex landscapes. One example of a group optimization algorithm facing a rugged landscape is the Particle Swarm Optimization (PSO) algorithm. When facing rugged landscapes with many local optima, PSO often gets trapped in these local optima and fails to find the global optimum[27].
@@ 3.2 Position Update
As for random factors $r_{1}$ and $r_{2}$ shown in equation (4) and equation (5), $r_{1}$ is for controlling the magnitude of an escape step, making HEO escape behaviors more efficient than simple grid search.$r_{2}$ is added to randomize further the process. Both of those two random variables have exception 1 to ensure the feasibility. The random factor $r_{3}$ controls the global weight of this updating process, creating a fan area to search for the solution.
$\forall A,A=\{s_{1},s_{2}\},T(A)=s^{*},f(s^{*})=f^{*}$ (8)
$G_{P}=(S_{P},T),G_{O}=(S_{O},T),G_{H}=(S_{H},T)$ (9)
$s_{P}\in S_{P},s_{O}\in S_{O},s_{H}\in S_{O}$ (10)
Since the HEO originates from PSO, the paper could have a rough insight into its convergence from the aspect of its ancestor. Ignoring the random factors for adding randomness and the escape term contained in the formulas, the HEO has the same way of updating the position as PSO does.
Also, as equation (9) and (10) shown, due to the particles or quantum in the groups, the algebraic structure, being all restricted by the required bound, the group of HEO $G_{H}$, as well as $G_{P}$ in PSO, could both be technically defined as groups with a min function operation $T$ of equation (8). The individual that held local minimum cost is included in the individual’s history, which proves the boundness of it.
Secondly, the individual that has extreme value cost is in the span of the space, ensuring the invertibility of function $T$. Thirdly, as previously motioned, the HEO is the same as PSO without adding an ”escape term” as well as multiplying those factors, making them satisfy the commutative and associative law.
In the case of PSO as well as HEO, the identity element would be the particle or quantum with the lowest cost between two individuals(the local optimum and the current one), while the inverse identity of it would be the current individual, hence the identity as well as the inverse identity exists and unique. In conclusion, the algebraic structures $G_{P}$ and $G_{O}$ are groups.
$\because O(s_{P})=s_{H}$ (11)
$\therefore G_{H}=G_{P}\rtimes G_{O}$ (12)
Besides, the paper defined an operation or function that could convert element $s_{P}$ to $s_{H}$ in equation (11), and the group $G_{O}$ is a group of different values for scaling in HEO. Additionally, since the individuals are generated randomly, which means there a nearly zero odds for PSO and HEO to have overlap values, make the semi-direct product established [6].
The semidirect product is defined as the multiplication operation of two groups through Cartesian product, where the elements of the two groups are combined in a specific way to form a new group structure[6]. Consequently, the group $G_{H}$ in HEO can be viewed as the outcome of the semi-direct product between $G_{P}$ and the group of the additional term $G_{O}$ in equation (12), the extra term $G_{O}$ is converge due its exception, and the identity $e_{o}$ come from the identities of two groups. As per the definitions, the identity of the group is the local optimum that gets searched and decreases.
The established global convergence of Particle Swarm Optimization (PSO) [4] [21] implies, to a certain extent, the global convergence of HEO with an absence of the escape term ($c_{i}=0$). This conclusion is derived from the fact that the global optimum is achieved through the aggregation of local optima, thus demonstrating a comparable convergence behavior between the two algorithms.
Global convergence in swarm optimization algorithms is the ability of the entire swarm to reach the global optimal solution, as opposed to getting stuck in local optima. This is important because it ensures that the algorithm will find the best possible solution to the problem, which is critical for real-world applications where sub-optimal solutions can lead to subpar performance or failure. In other words, the global convergence of HEO also theoretically proves its effectiveness.
@@ 3.3 Vibration
Another behavior of HEO is common in other algorithms including AFSA, when the search is meaningless to making a random step for individual search. Yet those algorithms adopt the strategy, they ignore the influence of size of steps, and those search approaches highly rely on the hyper-parameters tunning. HEO tries to use the standard deviation of positions to judge the magnitude of one random step, creating a relatively stable performance compared to the others, this random search is called Vibration.
@@
Just like the behavior of quantum in physics, with higher energy level $a_{i}$, the quantum would be harder get influenced by others[24], which means smaller oscillations(seen in equation (13) and Fig.3), maintain an extent of diversity in the population. To maintain the convergence of the whole group, the increment of energy level $a_{i}$ only works in a few quantums. Additionally, other scaling approaches to adapting the space of high dimensions with gamma functions are computed much slower than just simply calculating the standard deviation of the position vectors. The reason for using Normal distribution due its universal applicability. According to the central limit theorem, with the increasing of the dimension, the size of the vibration is more approximate to the size of the step in the real situation. Since the HEO does not include periodical functions in this behavior, it cannot defined as oscillations, that’s the reason for its name.
@@ 3.4 Center Clipping
As Fig.4 illustrated, the quantum not only would be restricted by the searching area or constraints $S_{bound}$(the area with (a)) with half of side length $b$, but the group also be restricted by the $S_{group}$. $S_{group}$(the area of (b) in Fig.3)is the search space of the group that HEO allows to search with half of the side length $b_{g}$, the search space in HEO are all in the shape of a square, making it computing faster than calculating sphere. Moreover, as equation (18) shows, this bound of group $S_{group}$ is dependent on the differences between the current quantum’s position and the best one that HEO found. Additionally, this center clipping mechanism uses random factor $r_{5}$, making this bound soft edge and not converge too early and maintain the diversity of the group. The final result of the group gets into the range of $S_{result}$, which is the overlap area of the $S_{group}$ and $S_{bound}$.
@@ 3.5 Random Skip
As Fig.5 shows, the random skip behavior in HEO allows the entire group of quantum to explore a completely new and random search space, the blue area represents the search space and the orange part is for the constraint. This is achieved by randomly shifting each quantum’s position by a random vector $\vec{r}$ in equation (20). The components of the random vector $\vec{r}_{j}$ are uniformly sampled from the interval $[0,b]$, as shown in equation (22). There are numerous $b$ for $\vec{r}$, hence the shape of the search space also changed with it as in sub-figure(a)(b) does, also the space restricts the constraints as sub-figure(c) is shown. This random skip mechanism enables HEO to explore different regions of the search space and potentially discover new and better solutions.
@@ 4.2 Trend Analysis
For $F_{4}$, Fig.7 also shows the performances of HEO are much superior to GA, AFSA, and PSO. The HEO only gets around 1e-134 cost, and GWO achieves about 1e-50. The QPSO, AFSA, GWO, and GA express a similar trait on the plot, those algorithms all converge smoothly but stuck in optimums. Adversely, the oscillations of PSO might be brought by the complex landscape of the $F_{4}$ function, and the fluctuations of HEO’s plot might be caused by its unique random mechanism, such as random skip as well as its escape behavior.
@@
For $F_{11}$, Fig.11 shows HEO and GWO both get a reasonable cost around 1e-16, but do not get full convergence. The possible reason for explaining this result could be the hollows of the Ackley function $F_{11}$ in high dimension[3]. Those hollow landscapes with the same cost might be the issue that prevents further convergence of algorithms.
@@
For $F_{14}$, Fig.12 also shows the HEO found the best solution in the first few iterations. Besides, the GWO has a small inclination to converge further and possibly to find the global optima.
$\lim_{i\to\infty}x_{i+1}-x_{i}=0$ (23)
$\lim_{i\to\infty}E(v_{g})=\frac{(x_{g}-x_{i})}{2}=0\quad(c_{i}=0)$ (24)
$\lim_{i\to\infty}E(v_{l})=\frac{(x_{l}-x_{i})}{2}=0\quad(c_{i}=0)$ (25)
From analysis of 14 benchmark functions, illustrates Halfway Escape Optimization (HEO) gets successful convergences on each case, which validates with the assumption in equation (23). For an ideal situation, the paper could further validate this result by conducting from the equations (1) (2) and (3). When the escape value $c_{i}=0$, and $i$ approximate the infinity, which means the global optimum $x_{g}$ as well as $x_{l}$ should be equal to $x_{i}$, since $x_{i+1}=x_{i}$ shown in equation (23). Besides, the exception value of the random factor is all one except the factor $r_{3}$, which brings a 1/2 scaling.
$\lim_{i\to\infty}x_{i+1}-x_{i}=E(v_{l})+E(v_{h})=\frac{(x_{l}-x_{i})+(x_{g}-x_{i})}{2}=0$ (26)
In summary, as the (26) shows, the paper could summarize the ideal situation of the HEO’s exception value as equations (24) (25). By substituting these two terms back into equation (1), the paper can demonstrate the validity of this equation (23). In other words, the HEO could converge once it finds a better optimum,which proves its convergence in another aspect.
@@ 4.3 Experimental Comparison
As for consider about the efficiency, based on the data from Table 5, the search time of five algorithms in those functions always shows a relationship that: $T_{PSO}<T_{GA}<T_{HEO}\approx T_{QPSO}<T_{GWO}\approx T_{AFSA}$. In conclusion, the HEO demonstrates an acceptable search time as well as a good performance among test functions except for $F_{12}$, but further validation of its effectiveness is still needed. For example, for have a more comprehensive of view, further tests could focus on testing HEO in other objective functions and various dimensions instead of just 30. This broadened experiment could give a more precise evaluation of the HEO.
5 Traveling Salesman Problem (TSP)
The Traveling Salesman Problem (TSP) is a well-known optimization problem in computer science and operations research. It involves finding the shortest route for a salesman to visit a set of cities exactly once. TSP is an NP-hard problem, which means that as the problem size increases, the time complexity of solving it will significantly increase, making it a well case for testing HEO.
This study aims to evaluate the effectiveness of the HEO algorithm in solving general Path Planning Problems. To do so, the algorithm is compared against Tabu Search and Random Search, using randomly generated datasets with 25, 50, and 100 cities with 2-dimensional position ranges within [-10,10]. The performance of these algorithms is assessed based on 100 iterations of searching and starts with the first city.
Tabu Search is a popular heuristic algorithm widely used in combinatorial optimization problems[10]. It systematically explores the search space while avoiding local optima through the utilization of a ”tabu list” that prevents revisiting recently examined solutions. This mechanism allows for the exploration of new regions and the discovery of improved solutions. Tabu Search has gained prominence due to its efficacy in solving intricate optimization problems, making it a well-tested algorithm for comparison. Random Search is another usually used baseline for evaluation in TSP, which was first used as a benchmark in 1986 [11].
In Fig.13, the path taken by the algorithms with 25 cities is depicted for a clear overview. Additionally, Table 6 demonstrates that the HEO algorithm consistently outperforms Random Search in terms of achieving shorter distances across all test cases. As the number of cities increases, Tabu Search experiences a significant increase in execution time. For instance, when there are 25 cities, the Tabu Search only costs 1.334546 seconds for 100 iterations, which is a shorter time compared to HEO. But when the number of cities is larger than 25, the HEO shows a relatively shorter time for searches. This highlights the potential of HEO as an alternative for more complex scenarios.
Overall, this research highlights the effectiveness of the HEO algorithm in general Path Planning Problems, as verified through the comparison with Tabu Search and Random Search methodologies. The findings indicate that HEO offers acceptable performance and potential. Also, compared to specialized algorithms that specifically focus on local search, such as Simulated Annealing (SA) [16], HEO performs worse than Tabu Search. The Halfway Escape Optimization (HEO) has a relatively weaker capability for local search. However, the complexity of this test is relatively simple compared to the real-time applications.
Table 6: Search Time(s/1000 iters) of the Algorithms
Methods Cities Distance Search Time(s/100 iters)
Tabu 25 100.626530 1.334546
HEO 25 133.276518 3.973852
Random 25 198.440632 0.013264
Tabu 50 172.474598 6.105021
HEO 50 300.411801 3.707406
Random 50 447.679963 0.022514
Tabu 100 301.104570 88.637227
HEO 100 674.661410 10.675231
Random 100 923.988814 0.079656
[…]
Further validations of the HEO algorithm’s effectiveness in TSP can be conducted through more complex scenarios. For example, real-world problems, such as route planning for drone delivery, can be used to verify the algorithm’s capabilities [2]. Additionally, comparing HEO to other swarm optimization algorithms in form of TSP, can provide insights into the algorithm’s strengths and weaknesses in different contexts. Future research can also investigate the impact of different parameters, such as the scaling radius $R$ in (4), on the algorithm’s performance. Nonetheless, the results presented in this study demonstrate the potential of HEO as a promising algorithm for solving TSP and other path planning problems.
Figure 13: The Paths of the Methods in TSP
6 The Pressure Vessel Design Problem

### v4

@@ Abstract
This paper first proposes the Halfway Escape Optimization (HEO) algorithm, a quantum-inspired metaheuristic designed to address general optimization problems characterized by rugged landscapes and high-dimensionality with an efficient convergence rate. The study presents a comprehensive comparative evaluation of HEO’s performance against established optimization algorithms, including Particle Swarm Optimization (PSO), Genetic Algorithm (GA), Artificial Fish Swarm Algorithm (AFSA), Grey Wolf Optimizer (GWO), and Quantum behaved Particle Swarm Optimization (QPSO). The primary analysis encompasses 14 benchmark functions with dimension 30, demonstrating HEO’s effectiveness and adaptability in navigating general optimization landscapes and providing valuable insights into its performance. The test of HEO in Pressure Vessel Design and Tubular Column Design infers its feasibility and potential in real-time applications. Further validation in Osmancik-97 and Cammeo Rice Classification proves the effectiveness of HEO and achieves a higher accuracy record.
@@
we propose a new algorithm called HEO we provide a comprehensive benchmark of 14 Objective Functions for our HEO algorithm we further evaluate the effectiveness of HEO in 3 classical engineering problems we optimize Logistic Regression model with HEO, achieving a higher record in Osmancik-97 and Cammeo Rice Classification
@@ 1 Introduction
The main motivation behind the development of the Halfway Escape Optimization (HEO) algorithm stems from the limitations of existing optimization methods in terms of efficiency and adaptability to various single-objective optimization problems. HEO is a quantum-inspired metaheuristic designed to tackle the complexities of rugged landscapes, high-dimensionality, and multimodal search spaces with a focus on achieving fast convergence rates and reducing the time cost in searching. Inspired by the behavior of quantum particles and the concept of halfway escape, the algorithm offers a versatile and adaptive approach to exploration and exploitation in challenging optimization domains.
This paper provides a comprehensive analysis and evaluation of the HEO algorithm, focusing on its core principles, adaptability, and robustness in navigating diverse optimization landscapes. The unique energy-driven behavior, vibration strategies, and exploratory mechanisms of HEO are examined in the context of solving a range of simple as well as complex benchmark functions. The evaluations demonstrate HEO’s effectiveness in achieving convergence, balancing exploration and exploitation, and discovering high-quality solutions across a variety of challenging optimization problems.
Comparative studies with other swarm optimization algorithms highlight HEO’s adaptability and robustness in addressing the complexities of multimodal and high-dimensional search spaces, showcasing its fast convergence speed. The results of this study underscore the importance of the HEO algorithm as a promising approach to address complex optimization challenges, with implications for a wide range of practical applications, including structural design[7][8], path planning[21], circuit planning[18], or model optimization[23].
The detailed analysis presented demonstrates the potential of the HEO algorithm as a powerful and adaptive optimization method. Overall, the paper first proposes HEO method and primary testing its effectiveness and convergence in 14 single objective functions.
@@ 2 Related Work
In the realm of optimization, metaheuristic algorithms play a key role in the addressing of challenging optimization problems. These algorithms draw inspiration from natural phenomena and collective behaviors to navigate high-dimensional and multimodal search spaces, offering adaptive and robust solutions.
@@
Particle Swarm Optimization (PSO) has gained widespread recognition as a bio-inspired metaheuristic algorithm rooted in the collective behavior of bird flocks. PSO was first proposed by Eberhart and Kennedy in 1995. PSO iteratively updates the positions and velocities of the particles based on their best individual and global solutions, enabling effective exploration and exploitation of the search space[6]. While PSO has exhibited robustness in addressing various optimization problems, its potential for premature convergence and challenges in handling complex multimodal landscapes have been noted[22].
Ant Colony Optimization (ACO) is an algorithm inspired by the foraging behavior of ants. It was initially proposed by Marco Dorigo in his doctoral thesis in 1992 and has since gained significant attention in the field of optimization[5]. ACO simulates the behavior of ants in finding the shortest path between their nest and food source by depositing pheromone trails. These trails act as a form of communication, allowing ants to navigate and explore the search space effectively. ACO has been successfully applied to a wide range of optimization problems, including the well-known Traveling Salesman Problem (TSP)[21] and Vehicle Routing Problem (VRP)[5]. Its ability to find near-optimal solutions and its adaptability to dynamic environments make ACO a valuable algorithm for solving complex optimization problems.
Differential Evolution (DE) is another population-based stochastic optimization algorithm that was introduced by [20]. DE operates by iteratively evolving a population of candidate solutions through a combination of mutation, crossover, and selection operations. It is inspired by the process of natural evolution and survival of the fittest. DE utilizes the differential operators to explore and exploit the search space efficiently. It has been widely applied to various optimization problems and has shown promising performance in solving complex and multimodal functions. The simplicity, effectiveness, and robustness of DE make it a popular choice for solving real-world optimization problems.
@@
These existing metaheuristic algorithms have contributed significantly to the field of optimization, each with its unique strengths and limitations, the development of the Halfway Escape Optimization (HEO) algorithm seeks to offer an adaptive approach to address the general challenges including rugged landscapes, high-dimensionality, and multimodal search spaces [22].
To overcome the challenges presented by rugged landscapes, high-dimensionality, and multimodal search spaces, a new stable approach is required. Yet QPSO and HEO are both quantum-inspired algorithms with the same mechanism in computing global weight for updating position(seen in equation (6)), they still have significant differences. In contrast to the Quantum Particle Swarm Optimization (QPSO), the Halfway Escape Optimization (HEO) algorithm aims to provide such an approach by leveraging insights from quantum patterns and halfway escape strategies.
@@ 3 Proposed Methodology
In recent years, swarm optimization algorithms have been widely applied in various fields, such as machine learning, engineering design. However, these algorithms often face difficulties when faced with complex landscapes. One example of a group optimization algorithm facing a rugged landscape is the Particle Swarm Optimization (PSO) algorithm. When facing rugged landscapes with many local optima, PSO often gets trapped in these local optima and fails to find the global optimum[22].
@@ 3.2 Position Update
As for the random factors $r_{1}$ and $r_{2}$ shown in equation (4) and equation (5), $r_{1}$ is used to control the magnitude of an escape step, making the HEO escape behaviors more efficient than the simple grid search.$r_{2}$ is added to further randomize the process. Both of those two random variables have exception 1 to ensure the feasibility. The random factor $r_{3}$ controls the global weight of this update process, creating a fan area to search for the solution.
@@ 3.3 Vibration
Another behavior of HEO is common in other algorithms, including AFSA, when the search is meaningless to make a random step for an individual search. Yet those algorithms adopt the strategy, they ignore the influence of size of steps, and those search approaches highly rely on the hyper-parameters tunning. HEO tries to use the standard deviation of positions to judge the magnitude of one random step, creating a relatively stable performance compared to the others, this random search is called Vibration.
@@
Just like the behavior of quantum in physics, with higher energy level $a_{i}$, the quantum would be harder get influenced by others[19], which means smaller oscillations(seen in equation (8) and Fig.3), maintain an extent of diversity in the population. To maintain the convergence of the whole group, the increment of energy level $a_{i}$ only works in a few quantums. Additionally, other scaling approaches to adapting the space of high dimensions with gamma functions are computed much slower than just simply calculating the standard deviation of the position vectors. The reason for using Normal distribution due its universal applicability. According to the central limit theorem, with the increasing of the dimension, the size of the vibration is more approximate to the size of the step in the real situation. Since the HEO does not include periodical functions in this behavior, it cannot be defined as oscillations, that’s the reason for its name.
@@ 3.4 Center Clipping
As Fig.4 illustrated, the quantum not only would be restricted by the search area or constraints $S_{bound}$(the area with (a)) with half of the side length $b$, but the group also would be restricted by $S_{group}$. $S_{group}$(the area of (b) in Fig.3)is the search space of the group that HEO allows to search with half of the side length $b_{g}$, the search space in HEO are all in the shape of a square, making it computing faster than calculating sphere. Moreover, as equation (13) shows, this bound of group $S_{group}$ depends on the differences between the current quantum’s position and the best one that HEO found. Additionally, this center clipping mechanism uses a random factor $r_{5}$, which makes this bound soft edge not converge too early and maintains the diversity of the group. The final result of the group gets into the range of $S_{result}$, which is the overlap area of the $S_{group}$ and $S_{bound}$.
@@ 3.5 Random Skip
As Fig.5 shows, the random skip behavior in HEO allows the entire group of quantum to explore a completely new and random search space, the blue area represents the search space, and the orange part is for the constraint. This is achieved by randomly shifting the position of each quantum by a random vector $\vec{r}$ in Equation (15). The components of the random vector $\vec{r}_{j}$ are uniformly sampled from the interval $[0,b]$, as shown in equation (17). There are numerous $b$ for $\vec{r}$, hence the shape of the search space also changed with it as in sub-figure(a)(b) does, also the space restricts the constraints as sub-figure(c) is shown. This random skip mechanism enables HEO to explore different regions of the search space and potentially discover new and better solutions.
@@ 4.2 Trend Analysis
For $F_{4}$, Fig.7 also shows that the performances of HEO are much superior to GA, AFSA, and PSO. The HEO only gets around 1e-134 cost, and GWO achieves about 1e-50. The QPSO, AFSA, GWO, and GA express a similar trait on the plot, those algorithms all converge smoothly but stuck in optimums. Adversely, the oscillations of PSO could be caused by the complex landscape of the $F_{4}$ function, and the fluctuations of the HEO plot could be caused by its unique random mechanism, such as random skip and its escape behavior.
@@
For $F_{11}$, Fig.11 shows that HEO and GWO both obtain a reasonable cost around 1e-16, but do not get full convergence. The possible reason for explaining this result could be the hollows of the Ackley function $F_{11}$ in high dimension[2]. Those hollow landscapes with the same cost might be the issue that prevents further convergence of algorithms.
@@
For $F_{14}$, Fig.12 also shows the HEO found the best solution in the first few iterations. In addition, the GWO has a small inclination to converge further and possibly to find the global optima.
In summary, from analysis of 14 benchmark functions, illustrates Halfway Escape Optimization (HEO) gets successful convergences on each case, which validates with the effectiveness of HEO.
@@ 4.3 Experimental Comparison
For the efficiency, based on the data from Table 5, the search time of five algorithms in these functions always shows a relationship that: $T_{PSO}<T_{GA}<T_{HEO}\approx T_{QPSO}<T_{GWO}\approx T_{AFSA}$. In conclusion, the HEO demonstrates an acceptable search time as well as a good performance among test functions except for $F_{12}$, but further validation of its effectiveness is still needed. For example, for a more comprehensive view, further tests could focus on testing HEO in other objective functions and various dimensions instead of just 30. This broadened experiment could give a more precise evaluation of the HEO.
5 Stimulated Engineering Problems
5.1 The Pressure Vessel Design Problem